# On-device Sora: Enabling Diffusion-Based Text-to-Video Generation for Mobile Devices

### v1

@@ 5.2. Temporal Dimension Token Merging
Temporal Dimension Token Merging can be selectively applied during the denoising process to minimize potential negative impacts on video quality that may arise from processing merged tokens. Specifically, out of a total of $K$ denoising steps, tokens can be merged only for the initial $k$ steps, while the tokens for the remaining $K{-}k$ steps remain unmerged. This is based on the observation that, when tokens are merged along the temporal dimension, the noise values vary slightly across frames—a phenomenon not observed in image diffusion (Saharia et al., 2022) that does not involve a temporal dimension in the generation process. However, because the noise values in the early denoising steps are less subtle and critical, it is expected that applying token merging exclusively for initial steps does not substantially drop video quality.
6. Concurrent Inference with Dynamic Loading
@@
We tackle the challenge of limited device memory in text-to-video generation, restricting the on-device inference of large diffusion-based models, by introducing Concurrent Inference with Dynamic Loading (CI-DL), which partitions models and executes them in a concurrent and dynamic manner.
6.1. Concurrent Inference
@@
The model components of On-device Sora (Liu et al., 2024), i.e., STDiT (Zheng et al., 2024) and T5 (Raffel et al., 2020), easily exceed the available memory of many mobile devices, e.g., 3.3 GB RAM of iPhone 15 Pro, as shown in Figure 3. Given that the Transformer architecture (Wolf et al., 2020), which is the backbone for both T5 (Raffel et al., 2020) and STDiT (Zheng et al., 2024), we partition these models into smaller blocks (segments) and load them into memory accordingly for model inference.
Figure 8. The block loading and inference cycles for (a) sequential loading and inference, and (b) concurrent inference.
@@
To execute model inference using the model partitioning, each block must be loaded sequentially before execution, increasing the overall latency of video generation by incurring block loading time. Figure 8-(a) shows the sequential block load and inference cycle of STDiT (Zheng et al., 2024), where GPU remains idle intermittently, waiting for each block to be loaded into memory, and only begins execution after the loading is complete. This sequential loading and execution process significantly increases the overall latency of model inference.
@@
To minimize the increase in model inference latency caused by sequential block loading and execution, we propose Concurrent Inference, which leverages both the CPU and GPU for parallel block loading and execution; CPU loads the $(i+1)$th block, while GPU concurrently executes the $i$th block. Initially, the first and second blocks are loaded into memory concurrently, with the first block completing its loading first. Subsequently, the inference of the first block and the loading of the second block occur simultaneously. This process continues such that the inference of the $i$th block and the loading of the $(i+1)$th block overlap, ensuring continuous parallelism until the final block. Figure 8-(b) depicts the load and inference cycle of STDiT with Concurrent Inference, which shows that GPU is active without almost no idle time by performing block loading and inference in parallel.
@@
Given the large number of denoising steps performed by STDiT, which is partitioned into multiple blocks for execution, similar to T5 (Raffel et al., 2020), the number of blocks $b$ is expected to be large, leading to a significant reduction in latency. It is expected to accelerate the overall model inference effectively regardless of the device’s memory capacity. When the available memory is limited, then $b$ increases, while with larger memory, both $l$ and $e$ increase. In either case, it can result in an almost constant latency reduction $r$ in Equation 10.
Figure 9. Example videos generated by On-device Sora and Open-Sora (Zheng et al., 2024) (68 frames, 256×256 resolution).
6.2. Dynamic Loading
To further enhance the model inference latency, we propose Dynamic Loading, which is applied in conjunction with Concurrent Inference. It maintains a subset of model blocks in memory without unloading them, with the selection for the subset of blocks to be retained in memory dynamically determined based on the device’s available memory at runtime.
@@ 6.2. Dynamic Loading
Figure 10 depicts Dynamic Loading; the first four model blocks are loaded in a retrained state. Unlike other blocks, e.g., the 5th block, these blocks are not unloaded to memory after the initial step, reducing block loading overhead.
Figure 10. The block loading and inference cycle for Dynamic Loading applied with Concurrent Inference.
Applying Dynamic Loading, the latency reduction $r$ in Equation 10 for Concurrent Inference is updated as:
(11) $r=(b-d)\cdot\min(l,e)+d\cdot l-\alpha\cdot(1-d/b)$
where $d$ is the number of blocks maintained in memory. As the number of model blocks retained in memory increases (i.e., a larger $d$), the overhead of loading unloading blocks is further minimized, fully accelerating the overall model inference under the device’s run-time memory constraints. Dynamically keeping some model blocks in memory with a retrained state is particularly advantageous for STDiT (Zheng et al., 2024) that is iteratively executed for denoising steps, which entails loading and reusing the same blocks for each step.
7. Implementation
We implement On-device Sora on iPhone 15 Pro (Apple, 2023), leveraging its GPU of 2.15 TFLOPS and 3.3 GB of available memory, with the three methods proposed in Sec. 4, 5, and 6. The model components—T5 (Raffel et al., 2020), STDiT (Zheng et al., 2024), and VAE (Doersch, 2016)—in PyTorch (Paszke et al., 2019) are converted to MLPackage, an Apple’s CoreML framework (Sahin and Sahin, 2021) for machine learning apps. Since some state-of-the-art diffusion-related operations required for text-to-video generation are not supported by the current version of CoreML (Apple, 2023), we develop custom solutions, such as xFormer (Lefaudeux et al., 2022) and cache-based acceleration for model inference. Additional implementations, such as denoising scheduling, sampling pipeline, and tensor-to-video conversion, are developed in Swift (Apple, 2024) by utilizing Apple-provided and authenticated libraries. To optimize the models, T5 (Raffel et al., 2020), the largest model used in video generation, is quantized to int8, while other models operated in float32. We found that quantizing STDiT (Zheng et al., 2024) and VAE (Doersch, 2016) is challenging due to their sensitivity and performance degradation in video generation. Our future implementations of On-device Sora will explore additional optimization to further enhance model efficiency.
8. Experiment
Table 2. The VBench (Huang et al., 2024) evaluation by category: On-device Sora vs. Open-Sora (Zheng et al., 2024) (68 frames, 256×256 resolution).
@@ 8. Experiment
Figure 11. A visual comparison of videos generated by On-device Sora and Open-Sora (Zheng et al., 2024), evaluated using VBench (Huang et al., 2024).
8.1. Video Generation Performance
We evaluate the quality of videos generated on an iPhone 15 Pro (Apple, 2023), in comparison to videos produced by Open-Sora (Zheng et al., 2024) running on NVIDIA A6000 GPUs. To assess both temporal and frame-level video quality, we utilize VBench (Huang et al., 2024), the state-of-the-art benchmark for text-to-video generation, which provides a comprehensive evaluation metrics, including subject consistency, background consistency, temporal flickering, motion smoothness, dynamic degree, aesthetic quality, and imaging quality. The evaluation is conducted on 68-frame videos at 256×256 resolution, using text prompts provided by VBench (Huang et al., 2024), consisting of 100 examples each across eight categories: animals, architecture, food, humans, lifestyle, plants, scenery, and vehicles. Table 2 and Figure 11 present the comparison of the generated videos based on those metrics. The results demonstrate that On-device Sora generates videos with quality nearly equivalent to Open-Sora in most evaluation metrics, exhibiting only a slight drop in frame-level quality, averaging 0.029, while achieving an average improvement of 0.07 in dynamic degree.
Figure 9 shows example prompts and videos, compared with Open-Sora (Zheng et al., 2024). For the prompt ”a stack of dried leaves burning in a forest”, both On-device Sora and Open-Sora generate visually plausible videos, both illustrating stack of burning dry leaves and forest in the background. Similarly, for ”close-up of a lemur”, both models produce descriptive videos: On-device Sora shows a lemur turning its head, while Open-Sora delivers a less dynamic yet visually comparable depiction.
Table 3. The video quality and generation speedup under different settings of LPL (Linear Proportional Leap).
[…]
Subject
Consistency
Background
Consistency
Temporal
Flickering
Motion
Smoothness
Dynamic
Degree
Aesthetic
Quality
Imaging
Quality
16/30 (53%) 0.805 527.27 0.97 0.97 0.99 0.99 0.18 0.50 0.55 1.94$\times$
21/30 (70%) 0.832 344.40 0.97 0.97 0.99 0.99 0.20 0.50 0.56 1.49$\times$
23/30 (76%) 0.840 305.47 0.97 0.97 0.99 0.99 0.21 0.50 0.56 1.34$\times$
25/30 (80%) 0.848 276.68 0.97 0.97 0.99 0.99 0.21 0.50 0.56 1.24$\times$
30/30 (100%) - - 0.97 0.97 0.99 0.99 0.21 0.50 0.57 1.00$\times$
Dynamic ($\mu$:17.73/30) 0.827 370.86 0.97 0.97 0.99 0.99 0.20 0.50 0.56 1.53$\times$
Table 4. The video quality and speedup under different merging steps of TDTM (Temporal Dimension Token Merging).
[…]
Subject
Consistency
Background
Consistency
Temporal
Flickering
Motion
Smoothness
Dynamic
Degree
Aesthetic
Quality
Imaging
Quality
30/30 (100%) 0.595 1225.69 0.97 0.97 0.99 0.99 0.06 0.50 0.56 1.27$\times$
25/30 (83%) 0.599 1168.85 0.97 0.97 0.99 0.99 0.06 0.50 0.56 1.21$\times$
20/30 (66%) 0.604 1056.47 0.97 0.97 0.99 0.99 0.06 0.50 0.56 1.16$\times$
15/30 (50%) 0.612 924.92 0.97 0.97 0.99 0.99 0.12 0.50 0.57 1.13$\times$
10/30 (33%) 0.622 784.67 0.97 0.97 0.99 0.99 0.16 0.50 0.56 1.10$\times$
0/30 (0%) - - 0.96 0.97 0.99 0.99 0.23 0.50 0.58 1.00$\times$
8.2. Linear Proportional Leap
Table 3 presents the video generation performance and speed-up of On-device Sora when applying Linear Proportional Leap (LPL) in Sec. 4. In the table, ‘LPL Setting’ indicates the number of denoising steps used for video generation out of a total of 30 steps, while the remaining steps are omitted by LPL. We also evaluate a dynamic version of LPL, referred to as ‘Dynamic’ in Table 3, which determines the number of denoising steps at runtime based on the cosine similarities between two adjacent drifts estimated using STDiT (Zheng et al., 2024). The dynamic LPL stops denoising and makes a proportional leap toward the final video when three cosine similarity measurements fail to improve with tolerance of $10^{-4}$, with a minimum 15 denoising steps (50%). Overall, LPL enable the generation of videos whose quality, as measured by SSIM (Fan et al., 2019), FVD (Unterthiner et al., 2019), and VBench (Huang et al., 2024), is comparable to that of Open-Sora (Zheng et al., 2024), without noticeable visual degradation (e.g., 0.740 vs. 0.743 in average of VBench (Huang et al., 2024)), while accelerating video generation up to 1.94$\times$. It is shown that LPL can reduce denoising latency linearly without incurring computational overhead. By directly reducing denoising steps, it facilitates efficient video generation while maintaining robust performance.
Figure 12 presents example videos generated using LPL, where all LPL settings consistently produce semantically identical target videos, with most video quality remaining stable across various numbers of denoising steps.
Figure 12. The snapshots of videos (68 frames, 256×256 resolution) applied with various LPL settings (Table 3).
8.3. Temporal Dimension Token Merging
Table 4 presents the video quality and video generation speedup achieved when varying numbers of denoising steps to which Temporal Dimension Token Merging (TDTM) (Sec. 5) is applied, indicated as ‘Merging Steps’ in the table, out of a total of 30 denoising steps. Each configuration of merging steps is evaluated on 68-frame videos at a resolution of 256×256 pixels, using SSIM (Fan et al., 2019), FVD (Unterthiner et al., 2019), and VBench (Huang et al., 2024) as evaluation metrics. The results indicate that increasing the number of merging steps consistently accelerates video generation, ranging from 1.10×1.10$\times$1.10 × to 1.27×1.27$\times$1.27 × speedups, while maintaining stable quality metrics; the average scores for SSIM, FVD, and VBench remain at 0.606, 1032.120, and 0.725, respectively. Nonetheless, declines in the dynamic degree metric is observed, revealing a trade-off between maintaining visual dynamics and reducing token processing complexity. This indicates the importance of striking a balance between video dynamics and speedup. We found that selectively applying TDTM to specific denoising steps can effectively reduce visual noises. For instance, limiting TDTM to the first 15 denoising steps while not applying it to the rest steps tends to result in a less severe quality drop compared to applying TDTM to all steps, which can mitigate issues like flickering or dynamic degree to provide improved video quality.
Figure 13. The snapshots of videos (68 frames, 256×256 resolution) applied with various merging steps of TDTM (Table 4).
Figure 13 shows examples of video frames generated with varying numbers of denoising steps to which TDTM is applied, out of 30 steps. The results demonstrate that even as TDTM is applied to an increasing number of denoising steps, the quality of the video frames seems to remain consistent.
@@ 8.3. Temporal Dimension Token Merging
8.4. Concurrent Inference with Dynamic Loading

### v2

@@ 5.2 Temporal Dimension Token Merging
Temporal Dimension Token Merging can be selectively applied during the denoising process to minimize potential negative impacts on video quality that may arise from processing merged tokens. Specifically, out of a total of $K$ denoising steps, tokens can be merged only for the initial $k$ steps, while the tokens for the remaining $K{-}k$ steps remain unmerged. This is based on the observation that, when tokens are merged along the temporal dimension, the noise values vary slightly across frames—a phenomenon not observed in image diffusion [53] that does not involve a temporal dimension in the generation process. However, because noise values in the early denoising steps are less critical, it is expected that applying token merging exclusively for initial steps does not substantially drop video quality.
6 Implementation
We implement the proposed On-device Sora on iPhone 15 Pro [1], leveraging its GPU of 2.15 TFLOPS and 3.3 GB of available memory, with the two methods proposed in Sec. 4 and 5. In addition, to execute large video generative models (i.e., T5 [48] and STDiT [83]) with the limited device memory, we devise and implement Concurrent Inference with Dynamic Loading (CI-DL), which partitions the models into smaller blocks that can be loaded into the memory and executed concurrently. The details of CI-DL is described in Sec. A.1. The model components—T5 [48], STDiT [83], and VAE [15]—in PyTorch [45] are converted to MLPackage, an Apple’s CoreML framework [54] for machine learning apps. Since current version of CoreML [1] lacks support for certain diffusion-related operations in text-to-video generation, we develop custom solutions like xFormer [33] and cache-based acceleration. We implement denoising scheduling, sampling pipeline, and tensor-to-video conversion in Swift [2] using Apple-provided libraries. To optimize models, T5 [48], the largest in video generation, is quantized to int8, while others models (STDiT [83] and VAE [15]) run in float32; we found that they are challenging to quantize due to sensitivity and performance degradation.
7 Experiment
We evaluate the performance of On-device Sora on 68-frame videos at 256×256 resolution, using 800 text prompts sampled from VBench [28] and 1,000 each from Pandas70M [11], and VidGen [61]. To assess both temporal and frame-level video quality, we utilize VBench [28], the state-of-the-art benchmark for text-to-video generation, which provides comprehensive metrics, including subject consistency, background consistency, temporal flickering, motion smoothness, dynamic degree, aesthetic quality, and imaging quality. Additional experiments on another text-to-video generation model, i.e., Pyramidal Flow [29], are provided in Fig. 15 and  Tab. 7 in App. B.
7.1 Video Generation Performance
Table 2: The VBench [28] evaluation summary: On-device Sora vs. Open-Sora [83] (68 frames, 256×256), full results in Tab. 6 in App. B.
[…]
      Subject
      Consistency
      Background
      Consistency
      Temporal
      Flickering
      Motion
      Smoothness
      Dynamic
      Degree
      Aesthetic
      Quality
      Imaging
      Quality
      Open-Sora       0.97       0.97       0.99       0.99       0.21       0.50       0.56
      On-device Sora       0.96       0.97       0.99       0.99       0.27       0.47       0.53
We evaluate the quality of videos generated by On-device Sora on the iPhone 15 Pro [1], in comparison to videos produced by Open-Sora [83] running on NVIDIA A6000 GPUs. Tab. 2 summarizes the averaged results that compare those videos evaluated with VBench [28], including categories such as animals, humans and lifestyle. Tab. 6 and Fig. 11 in App. B provide detailed full per-category results. The results demonstrate that On-device Sora generates videos with quality nearly equivalent to Open-Sora in most metrics, exhibiting only a slight drop in frame-level quality, averaging 0.03, while achieving an average 0.06 improvement in dynamic degree.
Fig. 6 shows example videos, compared with Open-Sora [83]. For the prompt “a stack of dried leaves burning in a forest”, both On-device Sora and Open-Sora generate visually plausible videos, illustrating stack of burning dry leaves and forest in the background. Similarly, for “close-up of a lemur”, both produce descriptive videos: On-device Sora shows a lemur turning its head, while Open-Sora delivers a less dynamic yet visually comparable depiction.
Table 3: The video quality and generation speedup under different settings of LPL (Linear Proportional Leap).
[…]
Subject
Consistency
Background
Consistency
Temporal
Flickering
Motion
Smoothness
Dynamic
Degree
Aesthetic
Quality
Imaging
Quality
VBench Dynamic ($\mu$:17.73/30) 0.97 0.97 0.99 0.99 0.20 0.50 0.56 1.53$\times$
16/30 (53%) 0.97 0.97 0.99 0.99 0.18 0.50 0.55 1.94$\times$
30/30 (100%) 0.97 0.97 0.99 0.99 0.21 0.50 0.57 1.00$\times$
Pandas70M Dynamic ($\mu$:17.66/30) 0.97 0.97 0.99 0.99 0.22 0.47 0.60 1.92$\times$
16/30 (53%) 0.97 0.97 0.99 0.99 0.19 0.47 0.60 1.62$\times$
30/30 (100%) 0.97 0.97 0.99 0.99 0.22 0.45 0.58 1.00$\times$
VidGen Dynamic ($\mu$:18.08/30) 0.95 0.96 0.98 0.99 0.36 0.47 0.57 1.94$\times$
16/30 (53%) 0.95 0.96 0.98 0.98 0.33 0.46 0.56 1.53$\times$
30/30 (100%) 0.96 0.97 0.99 0.99 0.39 0.48 0.57 1.00$\times$
Table 4: The video quality and speedup under different merging steps of TDTM (Temporal Dimension Token Merging).
[…]
Subject
Consistency
Background
Consistency
Temporal
Flickering
Motion
Smoothness
Dynamic
Degree
Aesthetic
Quality
Imaging
Quality
VBench 30/30 (100%) 0.97 0.97 0.99 0.99 0.06 0.50 0.56 1.27$\times$
15/30 (50%) 0.97 0.97 0.99 0.99 0.12 0.50 0.57 1.13$\times$
0/30 (0%) 0.96 0.97 0.99 0.99 0.23 0.50 0.58 1.00$\times$
Pandas70M 30/30 (100%) 0.98 0.97 0.99 0.99 0.09 0.48 0.59 1.23$\times$
15/30 (50%) 0.98 0.97 0.99 0.99 0.13 0.47 0.58 1.11$\times$
0/30 (0%) 0.97 0.97 0.99 0.99 0.22 0.45 0.58 1.00$\times$
VidGen 30/30 (100%) 0.97 0.97 0.99 0.99 0.14 0.50 0.59 1.70$\times$
15/30 (50%) 0.97 0.98 0.99 0.99 0.16 0.47 0.58 1.36$\times$
0/30 (0%) 0.96 0.97 0.99 0.99 0.39 0.48 0.57 1.00$\times$
7.2 Linear Proportional Leap
Tab. 3 presents the video generation performance and speedup of On-device Sora when applying Linear Proportional Leap (LPL) proposed in Sec. 4. In the table, ‘LPL Setting’ indicates the number of denoising steps used for video generation out of a total of 30 steps, while the remaining steps are omitted by LPL. We also evaluate a dynamic version of LPL, referred to as ‘Dynamic’ in Tab. 3, which determines the number of denoising steps at runtime based on the cosine similarities between two adjacent drifts estimated using STDiT [83]. The dynamic LPL halts denoising when a predefined number of cosine similarity measurements fail to improve beyond a $10^{-4}$ tolerance, after a designated minimum number of denoising steps (50%).
Overall, LPL enables video generation with quality comparable to Open-Sora [83], without noticeable visual degradation (e.g., 0.743 vs. 0.736 in average of VBench [28]), while accelerating video generation up to 1.94×1.94$\times$1.94 × without any model optimization or re-training. LPL reduces denoising latency linearly without extra computation, enabling efficient video generation while retaining robust performance.
Fig. 7 presents example videos generated using LPL, where all LPL settings consistently produce semantically identical target videos, with most video quality remaining stable across various numbers of denoising steps.
Figure 7: The snapshots of videos (68 frames, 256×256 resolution) applied with various LPL settings (Tab. 3).
7.3 Temporal Dimension Token Merging
Tab. 4 presents the video quality and video generation speedup achieved when varying numbers of denoising steps to which Temporal Dimension Token Merging (TDTM) (Sec. 5) is applied, indicated as ‘Merging Steps’ in the table, out of a total of 30 denoising steps. The results indicate that increasing the number of merging steps consistently accelerates video generation, ranging from $1\times1$ to 1.70×1.70$\times$1.70 × speedups, while maintaining stable quality metrics; the average scores for VBench remain at 0.736. Nonetheless, some declines in the dynamic degree metric are observed, revealing a trade-off between maintaining visual dynamics and reducing token processing complexity. This indicates the importance of striking a balance between video dynamics and speedup. We found that selectively applying TDTM to specific denoising steps can effectively reduce visual noises. For instance, limiting TDTM to the first 15 denoising steps while not applying it to the rest of the denoising steps tends to result in a less severe quality drop compared to applying TDTM to all steps, which can mitigate issues like flickering or dynamic degree to provide improved video quality.
Fig. 8 shows examples of video frames generated with varying numbers of denoising steps to which TDTM is applied, out of 30 steps. The results show that even as TDTM is applied to an increasing number of denoising steps, the quality of the video frames seems to remain consistent.
Figure 8: The snapshots of videos (68 frames, 256×256 resolution) applied with various merging steps of TDTM (Tab. 4).
7.4 Video Generation Latency
Tab. 5 shows video generation latencies of two resolutions when each of the proposed methodologies—LPL, TDTM, and CI-DL—is applied individually, as well as the latency when all of them are applied together (‘All’). Latencies are measured with LPL activated at the 15th denoising step and TDTM applied throughout all steps, reported as the mean of three independent experiments. ‘STDiT’ and ‘Total’ specify whether the latency is measured solely for STDiT [83] or for end-to-end video generation, including T5 [48] and VAE [15]. The results demonstrate substantial latency reductions for each methodology compared to the case without using the proposed methods, i.e., 293.51 vs. 1768.32 seconds (Tab. 1) for STDiT. For the 192$\times$192 resolution video, STDiT (denoising process) takes less than five minutes when all three methodologies are applied. Additionally, it indicates that the methodologies do not interfere with each other, instead work synergistically to enhance latency.
Table 5: Ablation study on video generation latency (s). ‘All’ denotes the combined application of LPL, TDTM, and CI-DL.
Resolution Measurement LPL TDTM CI-DL All
192$\times$192 STDiT 390.72 696.03 566.81 293.51
Total 514.24 823.03 691.25 416.50
256$\times$256 STDiT 573.88 965.40 947.67 454.48
Total 754.08 1148.63 1127.88 638.09
8 Related Work
On-device Video Generation. On-device video generation has recently gained attention due to the growing demand for real-time, efficient, and privacy-preserving content creation, and recently begun to be studied. MobileVD [71], based on UNet [51] in Stable Video Diffusion [4], reduces computational costs by employing pruning, including low-resolution fine-tuning, temporal multi-scaling, and optimizations in channel and time blocks. SnapGen-V [70] utilizes Stable Diffusion v1.5 [50] for processing image information while efficiently handling temporal information through Conv3D and attention-based layers. Furthermore, with fine-tuning, the denoising step is reduced. However, these approaches require substantial GPU resources for model optimization, e.g., MobileVD gone 100K training iterations on four A100 GPUs, and SnapGen-V utilized more than 150K iterations on 256 NVIDIA A100 GPUs. In contrast, On-Device Sora instantly applies without additional model training or optimization, removing the requirement for GPU resources.
Rectified Flow. While Open-Sora [83] reduces the number of denoising steps by leveraging Rectified Flow [38], most related approaches [16, 84] require conditioned model training and/or distillation [84]. In contrast, the proposed Linear Proportional Leap effectively reduces the denoising steps without a significant performance drop, as validated using VBench [28], without requiring model re-training or distillation. Notably, it can be easily activated at runtime by just calculating the cosine similarities of drifts between consecutive denoising steps.
Token Merging. Most token merging methods [6, 7, 17] primarily focus on image generation, where tokens are merged based on the spatial similarity rather than temporal similarity. Although some temporal token merging have been proposed, they are applied to models in other domains [18, 34, 9], not in video generation. As such, Temporal Dimension Token Merging is the first to apply token merging based on the successive similarities between frames in video generation. Additionally, while previous works apply token merging to self-attention due to performance degradation [6, 7, 17, 34], On-device Sora shows that token merging can be applied to cross-attention with minimal performance loss, achieving 50% merging ratio.
9 Conclusion
We propose On-device Sora, the first training-free solution for generating videos on mobile devices using diffusion-based models. It addresses key challenges in video generation to enable efficient on-device operation. The issue of extensive denoising steps is addressed through Linear Professional Leap, the challenge of handling large tokens is mitigated with Temporal Dimension Token Merging, and memory limitations are overcome through Concurrent Inference with Dynamic Loading. These proposed efficient on-device video generation methodologies are not limited to On-device Sora but are broadly applicable to various applications, providing advancements in on-device video generation without additional model re-training.
Appendix A Appendix - Implementation
A.1 Concurrent Inference with Dynamic Loading
@@
On-device Sora tackles the challenge of limited device memory in text-to-video generation, which restricts the on-device inference of large diffusion-based models, by introducing Concurrent Inference with Dynamic Loading (CI-DL), which partitions models and executes them in a concurrent and dynamic manner.
@@
Concurrent Inference The model components of On-device Sora [39], i.e., STDiT [83] and T5 [48], easily exceed the available memory of many mobile devices, e.g., 3.3 GB RAM of iPhone 15 Pro, as shown in Fig. 2. Given that the Transformer architecture [67], which is the backbone for both T5 [48] and STDiT [83], we partition these models into smaller blocks (segments) and load them into memory accordingly for model inference.
@@
To execute video model inference using the model partitioning, each block must be loaded onto memory sequentially before execution, increasing the overall latency of video generation by incurring block loading time. Fig. 9-(a) shows the sequential block load and inference cycle of STDiT [83], where GPU remains idle intermittently, waiting for each block to be loaded into memory, and only begins execution after the loading is complete. This sequential loading and execution process significantly increases the overall latency of model inference.
Figure 9: The block loading and inference cycles for (a) sequential loading and inference, and (b) concurrent inference.
@@
To minimize the increase in model inference latency caused by sequential block loading and execution, we propose Concurrent Inference, which leverages both the CPU and GPU for parallel block loading and execution; CPU loads the $(i+1)$th block, while GPU concurrently executes the $i$th block. Initially, the first and second blocks are loaded into memory concurrently, with the first block completing its loading first. Subsequently, the inference of the first block and the loading of the second block occur simultaneously. This process continues such that the inference of the $i$th block and the loading of the $(i+1)$th block overlap, ensuring continuous parallelism until the final block. Fig. 9-(b) depicts the load and inference cycle of STDiT with Concurrent Inference, which shows that GPU is active without almost no idle time by performing block loading and inference in parallel.
@@
Given the large number of denoising steps performed by STDiT, which is partitioned into multiple blocks for execution, similar to T5 [48], the number of blocks $b$ is expected to be large, leading to a significant reduction in latency. It is expected to accelerate the overall model inference effectively regardless of the device’s memory capacity. When the available memory is limited, then $b$ increases, while with larger memory, both $l$ and $e$ increase. In either case, it can result in an almost constant latency reduction $r$ in Eq. 10.
Dynamic Loading To further enhance the model inference latency, we propose Dynamic Loading, which is applied in conjunction with Concurrent Inference. It maintains a subset of model blocks in memory without unloading them, with the selection for the subset of blocks to be retained in memory dynamically determined based on the device’s available memory at runtime.
@@
Fig. 10 depicts Dynamic Loading; the first four model blocks are loaded in a retrained state. Unlike other blocks, e.g., the 5th block, these blocks are not unloaded to memory after the initial step, reducing block loading overhead.
Figure 10: The block loading and inference cycle for Dynamic Loading applied with Concurrent Inference.
Table 6: The VBench [28] evaluation by category: On-device Sora vs. Open-Sora [83] (68 frames, 256×256 resolution).
@@
Figure 11: A visual comparison of videos generated by On-device Sora and Open-Sora [83], evaluated using VBench [28].
@@
Applying Dynamic Loading, the latency reduction $r$ in Eq. 10 for Concurrent Inference is updated as:
$r=b\cdot\min(l,e)+d\cdot\max(0,l-e)-\alpha\cdot(1-d/b)$ (11)
where $d$ is the number of blocks maintained in memory. As the number of model blocks retained in memory increases (i.e., a larger $d$), the overhead of loading unloading blocks is further minimized, fully accelerating the overall model inference under the device’s run-time memory constraints. Dynamically keeping some model blocks in memory with a retrained state is particularly advantageous for STDiT [83] that is iteratively executed for denoising steps, which entails loading and reusing the same blocks for each step.
Appendix B Appendix - Additional Experiments
B.1 Video Generation Performance
In Sec. 7.1, we evaluate the quality of videos generated on an iPhone 15 Pro [1] and compare them to videos produced by Open-Sora running on NVIDIA A6000 GPUs using VBench [28] as the evaluation benchmark. The evaluation is conducted on 68-frame videos at 256×256 resolution, using text prompts provided by VBench [28], consisting of 100 examples each across eight categories: animals, architecture, food, humans, lifestyle, plants, scenery, and vehicles. Tab. 6 and Fig. 11 present the comprehensive comparison of generated videos on all categories.
B.2 Concurrent Inference with Dynamic Loading